# Two-Timescale Design for Movable Antennas Enabled-Multiuser MIMO Systems

### v1

@@ V Simulation Results
    FPA with optimal beamforming: “FPA-OPT (opt $\boldsymbol{W}$)” applies optimal adaptive beamforming to the FPA system, using instantaneous CSI for performance comparison.
@@
Figs. 5 and 6 evaluate the ergodic sum rate performance of various schemes across different transmit power levels with $\kappa=100$ and $\kappa=6$, respectively. In both figures, MA-ZF consistently outperforms the other approaches across all transmit power levels except for the very low power scenario $P_{\text{tot}}=9dBm$, effectively exploiting spatial degrees of freedom provided by antenna repositioning. In contrast, MA-MRT exhibits a significantly lower ergodic sum rate compared to MA-ZF, highlighting the limitations of MRT beamforming in managing multiuser interference. A key observation is the impact of the Rician factor on system performance. As the Rician factor decreases from $\kappa=100$ to $\kappa=6$ transitioning from strong to weaker LoS conditions, there is a noticeable decline in performance for MA-enabled schemes, particularly MA-MRT, due to the reduced availability of deterministic channel components. Encouragingly, MA-ZF maintains robustness across both strong and moderate LoS scenarios. In summary, MA-enabled systems exhibit superior or comparable performance to FPA-based systems, with MA-ZF consistently outperforming MA-MRT. These results highlight the performance gains achieved by optimizing antenna positions, especially in strong LoS environments. Nonetheless, it is important to note that MAs alone cannot replace well-designed transmit beamforming; instead, they work in synergy with beamforming to maximize overall system performance.
@@
Finally, Figs. 11 and 12 explore the relationship between the number of users $M$ and the ergodic sum rate for various MA and FPA configurations under ZF and MRT beamforming, with $N=6$ and $N=8$, respectively. From both figures, it is observed that MA-ZF consistently delivers performance gains as the number of users increases, maintaining superiority over other schemes, particularly in scenarios with more users. In contrast, MA-MRT shows a gradual increase in performance but struggles with multiuser interference, leading to a decline in performance as the number of users approaches the number of antennas. Under high Rician factor conditions, MA-MRT achieves sum rates comparable to FPA-ZF and notably outperforms FPA-MRT. Given its low complexity, MA-MRT offers a promising alternative to sophisticated beamforming designs in FPA systems. In low Rician factor environments, the overall performance trends remain similar, but the rates drop due to the reduced deterministic LoS component, with MA-MRT being particularly impacted, rendering it almost ineffective. Despite this, MA-ZF continues to deliver superior performance as the number of users increases, showcasing its robustness in both strong and moderate LoS conditions. Moreover, comparing the results for $N=6$ and $N=8$, the addition of antennas boosts sum rates across all configurations, with the improvement being more pronounced for MA-ZF. This highlights MA-ZF’s ability to leverage spatial diversity for effective multiuser interference management.
VI Conclusion
@@
Building upon the current two-timescale design, future research can further explore hierarchical or multi-timescale joint beamforming and antenna position optimization in MA-enabled systems. Specifically, by tailoring these adaptive optimizations to specific user densities, spatial distributions, and varying channel conditions or statistical characteristics, it is possible to develop more fine-grained, efficient, and practical system designs. These approaches could hold the potential to enhance the performance and practicality of the MA technique, thereby making MA-enabled MIMO systems a key solution for diverse next-generation wireless networks.

### v2

@@ VI Simulation Results
    FPA with optimal beamforming: “FPA-OPT (opt $\boldsymbol{W}$)” applies optimal adaptive beamforming to the FPA system, using instantaneous CSI for performance comparison [^7] In “FPA-OPT (opt W),” the beamforming vectors for a fixed antenna configuration are iteratively optimized to maximize the MU-MIMO sum rate by reformulating the non-convex problem through fractional programming with a quadratic transform.
@@
Figs. 5 and 6 evaluate the ergodic sum rate performance of various schemes across different transmit power levels with $\kappa=100$ and $\kappa=6$, respectively. In both figures, MA-ZF consistently outperforms the other approaches across all transmit power levels except for the very low power scenario $P_{\text{tot}}=9dBm$, effectively exploiting spatial degrees of freedom provided by antenna repositioning. In contrast, MA-MRT exhibits a significantly lower ergodic sum rate compared to MA-ZF, highlighting the limitations of MRT beamforming in managing multiuser interference. In high transmit power regimes (above 15 dBm) under weaker LoS conditions ($\kappa$=6), MA-MRT lags behind FPA-MRT mainly because the stronger random NLoS component diminishes gains from antenna repositioning and MA-MRT’s simplistic power allocation is less adaptive than FPA-MRT’s fractional programming approach. A key observation is the impact of the Rician factor on system performance. As the Rician factor decreases from $\kappa=100$ to $\kappa=6$ transitioning from strong to weaker LoS conditions, there is a noticeable decline in performance for MA-enabled schemes, particularly MA-MRT, due to the reduced availability of deterministic channel components. Encouragingly, MA-ZF maintains robustness across both strong and moderate LoS scenarios. In summary, MA-enabled systems exhibit superior or comparable performance to FPA-based systems, with MA-ZF consistently outperforming MA-MRT. These results highlight the performance gains achieved by optimizing antenna positions, especially in strong LoS environments. Nonetheless, it is important to note that MAs alone cannot replace well-designed transmit beamforming; instead, they work in synergy with beamforming to maximize overall system performance.
@@
Finally, Figs. 11 and 12 explore the relationship between the number of users MM and the ergodic sum rate for various MA and FPA configurations under ZF and MRT beamforming, with $N=6$ and $N=8$, respectively. From both figures, it is observed that MA-ZF consistently delivers performance gains as the number of users increases, maintaining superiority over other schemes, particularly in scenarios with more users. In contrast, MA-MRT shows a gradual increase in performance but struggles with multiuser interference, leading to a decline in performance as the number of users approaches the number of antennas. Under high Rician factor conditions, MA-MRT achieves sum rates comparable to FPA-ZF and notably outperforms FPA-MRT. Given its low complexity, MA-MRT offers a promising alternative to sophisticated beamforming designs in FPA systems. In low Rician factor environments, the overall performance trends remain similar, but the rates drop due to the reduced deterministic LoS component, with MA-MRT being particularly impacted, rendering it almost ineffective. Despite this, MA-ZF continues to deliver superior performance as the number of users increases, showcasing its robustness in both strong and moderate LoS conditions. Moreover, comparing the results for $N=6$ and $N=8$, the addition of antennas boosts sum rates across all configurations, with the improvement being more pronounced for MA-ZF. This highlights MA-ZF’s ability to leverage spatial diversity for effective multiuser interference management.
Figs. 13-16 evaluate the impact of different NLoS modeling approaches, that is, spatially correlated vs. simplified i.i.d., on the ergodic sum rate, where the i.i.d. assumption is imposed with a minimum antenna spacing of $\lambda/2$.
Figure 13: Ergodic sum rate evaluation for MA-MRT with different NLoS models ($N=8,M=5$).
Fig.13 illustrates the convergence of the ergodic sum rate for the two-timescale MA design under MRT beamforming. The green curve represents the NLoS component modeled as a spatially correlated Rayleigh process with a correlation matrix, while the brown curve assumes an i.i.d. complex Gaussian NLoS component with an identity covariance matrix, constrained by the minimum antenna spacing. The small final gap observed for Rician factors $\kappa=$ 5, 10, and 15 dB validates that the minimum spacing constraint effectively mitigates spatial correlation, justifying the use of the simplified NLoS model for computational efficiency.
Figure 14: Ergodic sum rate evaluation for MA-ZF with different NLoS Models ($N=8,M=5$).
Figure 15: Ergodic sum rate evaluation for MA-MRT with different NLoS Models ($N=6,M=5$).
Figure 16: Ergodic sum rate evaluation for MA-ZF with different NLoS Models ($N=6,M=5$).
Figs. 14 presents the ergodic sum rate evolution for an MA system with ZF beamforming. The blue curve accounts for spatial correlation in the NLoS component, while the red curve represents the i.i.d. assumption with an identity covariance matrix under the minimum spacing constraint. The small gap, within 1 bps/Hz, suggests that the i.i.d. model provides a reasonable approximation, particularly at higher Rician factors where the LoS component dominates. Figs. 15 and 16 further investigate the impact of spatial correlation using fewer transmit antennas. The results indicate that spatial correlation can slightly degrade or improve performance depending on antenna configuration and repositioning. Compared to MRT, ZF beamforming is more sensitive to spatial correlation effects. Nonetheless, the numerical results remain consistent with the simplified model under the minimum spacing constraint. Moreover, while spatial correlation effects intensify as the array size increases, the MA architecture inherently mitigates these effects. By leveraging spatial degrees of freedom without requiring additional antennas or RF chains, MAs provide a practical means to counteract correlation-induced performance variations.
Figs. 17 and 18 assess the sensitivity of the proposed two-timescale design to position estimation errors, evaluating the impact of inaccuracies in both user and antenna positions on the ergodic sum rate. User position errors are modeled as uniformly distributed within [1,1][-1,1] meters, while the antenna position errors ranged from $\left[-\frac{\lambda}{20},\frac{\lambda}{20}\right]$.
Figure 17: Performance degradation with user position error with $\kappa=20dB$.
Figure 18: Performance degradation with antenna position error with $\kappa=20dB$.
Figure 17 plots the ergodic sum rate as a function of user position errors (0 to 1 meter) under different configurations. The sharpest decline occurs within the first 0.1 meters, indicating high sensitivity to small errors. Figure 18 illustrates the degradation in ergodic sum rate due to antenna position errors (0% to 5% of $\lambda$), showing a similar sensitivity. The impact is particularly pronounced for the MA-MRT scheme, which relies heavily on precise antenna repositioning for optimal performance. In summary, user position errors at the decimeter scale induce phase mismatches in the channel matrix, affecting the LoS signal alignment between the BS and the user. Likewise, MA positioning inaccuracies pose a significant challenge, as mechanical movement errors can cause subwavelength deviations from the optimized results, leading to performance degradation. Despite these challenges, the numerical results suggest that our proposed optimization algorithms effectively identify local optima for the MA system, where small errors or deviations can cause significant performance loss [^8] To accommodate user mobility and avoid user positioning error, frequent updates of statistical CSI and the corresponding MA positions adjustment become necessary. In practice, the MA functionality can be selectively disabled or updated less frequently depending on the deployment scenario, making the performance gain optional when precise positioning cannot be maintained. In addition, the micro-electromechanical systems (MEMS)-integrated antenna and the liquid antenna are two alternative solutions for implementing MAs with very high positioning accuracy (micro- to even nano-meter scale) in portable devices with compact components..
Figs. 19 and 20 compare the ergodic sum rate achieved by our surrogate-based algorithms against a near-optimal brute-force solution. The brute-force approach searches for optimal antenna positions over a discrete grid with a predefined number of points. We evaluate two configurations: (i) $N=4,M=3$ within $x\in[-1.6\lambda,1.6\lambda],y\in[-1.6\lambda,1.6\lambda]$ and (ii) $N=6,M=5$ within $x\in[-1.6\lambda,1.6\lambda],y\in[-1.6\lambda,2.4\lambda]$, both under $\kappa=20dB$.
Figure 19: Performance comparison with the brute-force method with $N=4,M=3$.
Figure 20: Performance comparison with the brute-force method with $N=6,M=5$.
In Fig. 19, the dashed purple and orange lines represent brute-force results for MA-ZF and MA-MRT, respectively, while the solid lines correspond to our method. As grid resolution increases, the brute-force solution improves and stabilizes. Our method closely approximates brute-force performance, with only slight degradation. Notably, for an $N$-antenna MA array, the brute-force approach requires an $N$-dimensional exhaustive search over two spatial dimensions per antenna. With $D$ discrete points per axis, the total number of evaluations scales as $D^{2N}$, making it computationally prohibitive even for moderate $N$ (for example, $D=10,N=4$ requires $10^{8}$ evaluations. The trends in Fig. 20 mirror those in Fig. 19, with the surrogate-based method exhibiting a small performance gap within 6%, demonstrating its near-optimality. Although our approach sacrifices a marginal degree of optimality, it significantly reduces complexity in polynomial time, compared to the brute-force method on exponential scale, which becomes infeasible for practical MA systems.
VII Conclusion
@@
Building upon our current two-timescale design, future research will explore hierarchical or multi-timescale joint beamforming and antenna position optimization in MA-enabled systems, integrating techniques such as predictive CSI and reinforcement learning (RL) to dynamically adapt to fast-fading channels. Specifically, by tailoring these adaptive optimizations to user densities, spatial distributions, and varying channel conditions, the system can retain the benefits of large-scale statistical optimization and small-scale beamforming while incorporating flexible medium-timescale adjustments to handle significant channel variations. These approaches hold the potential to enhance the performance and practicality of the MA technique, balancing dynamic repositioning gains against practical limitations such as mechanical latency and power consumption, and thus position MA-enabled MIMO systems as a key solution for diverse next-generation wireless networks.